OpenStudy (jaredstone4):

Calculate the volume of S, where S is the solid which is bounded by the graph of z = x^2+1 and the planes z = 2y-1 and y = 2. I'm having trouble calculating the bounds of x. I believe that y goes from 0 to 2 and z goes from 0 to 2y-1, but I cannot figure out x. Please help!

OpenStudy (baru):

oh boy... this is hard to visualize... i am not totally sure, but i'll have a go

OpenStudy (baru):

the lower surface of the region is comprised of z=x^2 +1 and the upper by z=2y-1 so the limits for z are from 2y-1 to x^2 + 1 i feel that the projection of the required region on the x-y plane is a rectangle so the y bounds would be: y co-ordinate when the plane z=2y-1 hits z=x^2+1at its lowest point, which is z=1 =>y=1 and y=2 ( because of the vertical plane y=2)

OpenStudy (baru):

i.e y limits are from 1 to 2

OpenStudy (baru):

for the x limits, consider the parabola formed at the intersection of y=2 and z=x^2+1 the required x co-ordinates would be the furthest and nearest points along x, where the the plane z=2y-1 cuts this parabola i.e y=2 => z=2(2)-1 = x^2 +1 3=x^2 + 1 \(x=\pm \sqrt{2}\)

OpenStudy (baru):

edit: i have reversed the z limits

OpenStudy (baru):

ok... now i think projection on xy plane might be a parabola :/ @ganeshie8 help!

ganeshie8 (ganeshie8):

looks it is easier to setup the projection in xz plane...

ganeshie8 (ganeshie8):

https://www.wolframalpha.com/input/?i= \int_1^3+\int_%28-sqrt%28z-1%29%29^%28sqrt%28z-1%29%29+\int_%28%28z%2B1%29%2F2%29^2+1+dydxdz

ganeshie8 (ganeshie8):

\(\href{https:///www.wolframalpha.com/input/?i=\int_1^3+\int_%28-sqrt%28z-1%29%29^%28sqrt%28z-1%29%29+\int_%28%28z%2B1%29%2F2%29^2+1+dydxdz}{here}\)

OpenStudy (baru):

nice!